OpenStudy (anonymous):

y=-10/9x-4/5 in the lowest terms slope and y intercept

OpenStudy (aroub):

slope=-10/9 y-intercept=-4/5

OpenStudy (saifoo.khan):

Slope = -10/9 y-intercp = -4/5

OpenStudy (anonymous):

y = -10/9 x - 4/5 Slope is -10/9 = -1( 1/9) y intercept is -4/5